### v2

@@ 6.1. Q&A Interactions with CHOIR
Figure 7. Distribution of questions asked to CHOIR across four research labs during the one-month field study. The graph shows the number of questions submitted through public channels versus direct messages (DM), broken down by students (blue) and lab directors (green). L1 shows a distinctive pattern with predominantly public channel usage, while L2, L3, and L4 primarily used private DMs for Q&A interactions.:
@@ 6.1.2. How Lab Culture Shapes Question Asking
One notable pattern that we found from L1, as shown in Figure 7, was that most of the questioners asked CHOIR in a public channel instead of via DM, contrasting sharply with the other three labs. L1 students asked 21 questions publicly compared to only 5 private messages. Students’ motivation to ask public questions stemmed from altruistic and supportive attitudes to help others. S1-3 explained: “I think it can share to everybody about the answer[…] everybody can just check the Slack and see.” This visibility allowed students who had similar questions to benefit from others’ queries without needing to ask again. Some students also used CHOIR to support their lab. S1-4 stated, “I mostly used it to help [D1]. Add stuff that wasn’t already there.” The public nature of these interactions enabled collective awareness and knowledge sharing and also served as a catalyst for documentation improvements, with multiple students explicitly using CHOIR to identify gaps in the handbook.
@@ 6.3.2. Directors’ Update Practices and Member Engagement
Table 2. Documentation update patterns across research labs during the one-month field study, showing commits, files modified, and word changes for both direct GitHub updates and CHOIR-assisted updates.:
Lab #. Update Method Commits Files Words Added Words Deleted
L1 Direct Update 19 5 1153 63
CHOIR 0 0 0 0
L2 Direct Update 1 1 1 1
CHOIR 4 1 261 0
L3 Direct Update 0 0 0 0
CHOIR 1 1 109 1
L4 Direct Update 7 2 122 29
CHOIR 6 2 348 2
Subtotal Direct Update 27 8 1276 93
CHOIR 11 4 718 3
Total 38 12 1994 96
@@ 7.1. Privacy-Awareness Tension in AI-mediated Organizational Memory
Our findings revealed a fundamental tension between students’ privacy preferences and managers’ need for awareness in organizational memory systems. Because the majority of Q&A took place through direct messages to CHOIR without being shared, lab directors had limited visibility into knowledge gaps in the documentation. Directors expressed willingness to update information that their students could not access through interactions with CHOIR. This challenge reflects a broader issue in collaborative work: maintaining awareness of individual and group activities is essential for successful collaboration (Dourish and Bellotti, 1992; Tam and Greenberg, 2006), yet our study shows that AI-mediated knowledge retrieval can suffer from the same limitation.
@@ 7.2. Expanding the Perceived Boundary of Documentable Knowledge
Our study revealed a gap between the knowledge that lab members possessed and what they felt was appropriate to document. Many participants hesitated to share context-specific experiences that may be generalizable and beneficial to their peers. This reluctance appears to stem from an implicit expectation that documentation should be universal, authoritative, and permanent—creating a high bar for contribution that excludes valuable experiential knowledge and partial solutions. Participants also recognized value in case-based information and partial policies that could provide personalized guidance when comprehensive documentation was unavailable. This challenge is well-studied: from early work on organizational memory systems (Conklin, 1997) to research on knowledge sharing barriers (Ackerman et al., 2013), studies have consistently identified the difficulty of capturing informal knowledge due to cultural and technical barriers toward outcome-focused documentation. Prior work on enterprise wikis found that members only occasionally contributed despite recognizing documentation value, often requiring explicit requests from managers (Munson, 2008; Kiniti and Standing, 2013). CHOIR addressed this barrier by embedding documentation opportunities within natural conversation flows. Participants noted that the knowledge extraction feature made contributing informal knowledge feel less burdensome, as it transformed ongoing discussions into documentation without requiring separate writing effort. This suggests that conversational AI systems can provide lightweight pathways for capturing experiential knowledge that traditional documentation systems have struggled to elicit.
@@ 7.5. Limitations
Our study has several limitations that may impact the generalizability. First, our field deployment involved a small sample of participants (22 total) across four university research labs, all from STEM fields with high technical literacy. This limits generalizability to organizations with different technical backgrounds, hierarchical structures, or cultural norms around knowledge sharing. Second, the one-month deployment period was sufficient for observing initial adoption patterns but insufficient for evaluating the longer-term effectiveness of organizational memory systems. Our study focused on discovering potential effects rather than proving sustained efficacy.
@@
Fourth, one participating lab (L4) was led by the corresponding author, which may introduce potential bias despite our efforts to minimize researcher influence. While the corresponding author limited their role to managing the process, answering questions, and approving documents, and the first and second authors did not participate as study subjects, peer-pressure inherent in being in the same lab or the natural interest in a research project that is thematically similar to their own research could have influenced the L4 participants’ engagement patterns or feedback. However, the key findings and discussion points are valid as they are not disproportionately driven by L4 data alone.
Lastly, our study lacks a controlled baseline and quantitative performance measures beyond descriptive statistics, which limits our ability to make causal claims about CHOIR’s effectiveness compared to alternatives such as wiki pages or existing lab practices (e.g., Google Docs/Slack Pinned Messages). As a result, our findings are descriptive rather than comparative, and further research is needed to determine how specific features of CHOIR contribute to its impact.
@@ A.1. Question Answering Prompt
Think of yourself as a knowledgeable senior student or friendly
professor who’s always ready to help with questions.
I have access to the organization’s documentation and knowledge
base, so I can help you find information and provide guidance
based on what we have documented.
@@
IMPORTANT: First, determine if you can answer the user’s question
based ONLY on the provided references. Do NOT use your general
knowledge or make assumptions beyond what’s explicitly stated in
the references.
@@
1. "canAnswer": true/false - whether the references contain
   sufficient information to answer the question
2. "response": your answer based on the references, or a friendly
   explanation that you couldn’t find the information
@@
- Be friendly, approachable, and helpful - like a senior colleague
  who genuinely wants to help
- Answer ONLY based on the provided references below - do NOT add
  general knowledge
- If multiple references contain conflicting information,
  prioritize the first reference in the list
- If you cannot answer based on the references, encourage the user
  to ask others or start a discussion to help improve our
  documentation
- When you cannot fully answer a question, mention which related
  references you found and what information they contain
@@
Analyze whether you can answer based on the documentation and
provide your response as JSON.
@@ A.2. Knowledge Extraction Prompt
You are CHOIR, a documentation specialist. Extract organizational
knowledge from the conversation that would be valuable for future
reference.
Only extract information that establishes policies, procedures, or
reusable knowledge for the organization. Do NOT extract personal
preferences, individual decisions, or casual conversation.
Start with a descriptive markdown section title (# [Topic Name]),
then write the information in natural paragraph format. Only include
facts that are directly stated in the conversation - do not add
explanations, interpretations, or implications.
Do not include personal conversation details like who said what.
Always preserve any URLs mentioned.
@@
What information is shared in the conversation that should be
documented?
@@ A.3. Document Update Prompt
You are a document editor. Improve existing content by integrating
the provided knowledge.
@@
- PRIORITIZE updating/replacing existing content when knowledge
  provides better, more accurate, or more comprehensive information
- If knowledge contradicts existing content, prefer the knowledge
  (assume it’s more current/accurate)
- If knowledge complements existing content without contradiction,
  add it in matching format
- If knowledge provides more specific details about existing points,
  merge them into improved versions
@@
Approach: Update first, then add if needed. Create the most accurate
and comprehensive version.

### v3

@@ 6.1. Q&A Interactions with CHOIR
Figure 7. Distribution of questions asked to CHOIR across four research labs during the one-month field study. The graph shows the number of questions submitted through public channels versus direct messages (DM), broken down by students (blue) and lab directors (green). L1 shows a distinctive pattern with predominantly public channel usage, while L2, L3, and L4 primarily used private DMs for Q&A interactions.
A horizontal bar chart showing question counts across four labs. Each lab has two bars representing Channel and DM usage. L1 shows 14 student questions and 7 director questions in the channel, with 4 director DM questions. L2 shows 3 student questions in the channel and 30 student questions in DM. L3 shows 1 director channel question and 31 student DM questions. L4 shows 1 student public question and 16 student DM questions. The chart uses blue bars for students and green bars for lab directors, with directors shown on the left side of each stacked bar.
@@ 6.1.2. How Lab Culture Shapes Question Asking
One notable pattern that we found from L1, as shown in Figure 7, was that most of the questioners asked CHOIR in a public channel instead of via DM, contrasting sharply with the other three labs. L1 students asked 14 questions publicly, with none through private messages. Students’ motivation to ask public questions stemmed from altruistic and supportive attitudes to help others. S1-3 explained: “I think it can share to everybody about the answer[…] everybody can just check the Slack and see.” This visibility allowed students who had similar questions to benefit from others’ queries without needing to ask again. Some students also used CHOIR to support their lab. S1-4 stated, “I mostly used it to help [D1]. Add stuff that wasn’t already there.” The public nature of these interactions enabled collective awareness and knowledge sharing and also served as a catalyst for documentation improvements, with multiple students explicitly using CHOIR to identify gaps in the handbook.
@@ 6.3.1. Perceived Value of Document Update via CHOIR
Table 2. Documentation update patterns across research labs during the one-month field study, showing commits, files modified, and word changes for both direct GitHub updates and CHOIR-assisted updates.
A table with 5 columns showing documentation update statistics for 4 research labs (L1-L4). Each lab has two rows: Direct Update and CHOIR system updates. L1 shows 19 direct commits with 1153 words added and no CHOIR usage. L2 shows minimal direct updates (1 commit) but 4 CHOIR commits with 261 words added. L3 shows no direct updates and 1 CHOIR commit with 109 words added. L4 shows balanced usage with 7 direct commits and 6 CHOIR commits, totaling 470 words added across both methods.
Lab Update [1pt]Method Commits Files Words (+/-)
L1 Direct 19 5 +1153 / -63
CHOIR 0 0 0
L2 Direct 1 1 +1 / -1
CHOIR 4 1 +261 / -0
L3 Direct 0 0 0
CHOIR 1 1 +109 / -1
L4 Direct 7 2 +122 / -29
CHOIR 6 2 +348 / -2
Subtotal Direct 27 8 +1276 / -93
CHOIR 11 4 +718 / -3
Total 38 12 +1994 / -96
@@ 7.1. Privacy-Awareness Tension in AI-mediated Organizational Memory
Our findings revealed a fundamental tension between students’ privacy preferences and managers’ need for awareness in organizational memory systems. Because the majority of Q&A took place through direct messages to CHOIR without being shared, directors had limited visibility into knowledge gaps in the documentation. Directors expressed willingness to update information that their students could not access through interactions with CHOIR. This challenge reflects a broader issue in collaborative work: maintaining awareness of individual and group activities is essential for successful collaboration (Dourish and Bellotti, 1992; Tam and Greenberg, 2006), yet our study shows that AI-mediated knowledge retrieval can suffer from the same limitation.
@@ 7.2. Expanding the Perceived Boundary of Documentable Knowledge
Our study revealed a gap between the knowledge that lab members possessed and what they felt was appropriate to document. Many participants hesitated to share context-specific experiences that may be generalizable and beneficial to their peers. This reluctance appears to stem from an implicit expectation that documentation should be universal, authoritative, and permanent—creating a high bar for contribution that excludes valuable experiential knowledge and partial solutions. Participants also recognized value in case-based information and partial policies that could provide personalized guidance when comprehensive documentation was unavailable.
This challenge is well-studied: from early work on organizational memory systems (Conklin, 1997) to research on knowledge sharing barriers (Ackerman et al., 2013), studies have consistently identified the difficulty of capturing informal knowledge due to cultural and technical barriers toward outcome-focused documentation. Moreover, research has shown that knowledge needs in document work are inherently role-dependent and context-specific (Jahanbakhsh et al., 2022), challenging assumptions of universal documentation. Prior work on enterprise wikis found that members only occasionally contributed despite recognizing documentation value, often requiring explicit requests from managers (Munson, 2008; Kiniti and Standing, 2013). CHOIR addressed this barrier by embedding documentation opportunities within natural conversation flows. Participants noted that the knowledge extraction feature made contributing informal knowledge feel less burdensome, as it transformed ongoing discussions into documentation without requiring separate writing effort. This suggests that conversational AI systems can provide lightweight pathways for capturing experiential knowledge that traditional documentation systems have struggled to elicit.
@@ 7.5. Limitations
Our study has several limitations that may impact the generalizability. First, our field deployment involved a small sample of participants across four university research labs, all from STEM fields with high technical literacy. This limits generalizability to organizations with different technical backgrounds, hierarchical structures, or cultural norms around knowledge sharing. Second, the one-month deployment period was sufficient for observing initial adoption patterns but insufficient for evaluating the longer-term effectiveness of organizational memory systems. Our study focused on discovering potential effects rather than proving sustained efficacy.
@@
Fourth, one participating lab (L4) was led by the corresponding author, which may introduce potential bias despite our efforts to minimize researcher influence. L4 may also be particularly well aligned with the assumptions and practices that CHOIR supports, given the system’s origin in addressing needs observed in this context. While the corresponding author limited their role to managing the process, answering questions, and approving documents, and the first and second authors did not participate as study subjects, peer-pressure inherent in being in the same lab or the natural interest in a research project that is thematically similar to their own research could have influenced the L4 participants’ engagement patterns or feedback. However, the key findings and discussion points are valid as they are not disproportionately driven by L4 data alone.
Lastly, our study lacks a controlled baseline and quantitative performance measures beyond descriptive statistics, which limits our ability to make causal claims about CHOIR’s effectiveness compared to alternatives such as wiki pages or existing lab practices (e.g., Google Docs/Slack Pinned Messages). As a result, our findings are descriptive rather than comparative, and further research is needed to determine how specific features of CHOIR contribute to its impact. Future controlled experiments could compare an improved version of CHOIR against baseline conditions where shared documents and Slack are not connected through an AI mediator. Such studies could measure question response time, documentation update frequency and quality, knowledge accessibility, and participation in organizational memory maintenance.
@@ 8. Conclusion
9. Acknowledgments
We thank all professors and students who participated in the user studies, as well as the reviewers for their constructive feedback.
@@ A.1. Question Answering Prompt
Think of yourself as a knowledgeable senior student or friendly professor who’s always ready to help with questions.
I have access to the organization’s documentation and knowledge base,
so I can help you find information and provide guidance based on what we have documented.
@@
IMPORTANT: First, determine if you can answer the user’s question based ONLY on the provided references.
Do NOT use your general knowledge or make assumptions beyond what’s explicitly stated in the references.
@@
1. ‘‘canAnswer’’: true/false - whether the references contain sufficient information to answer the question
2. ‘‘response’’: your answer based on the references, or a friendly explanation that you couldn’t find the information
@@
- Be friendly, approachable, and helpful - like a senior colleague who genuinely wants to help
- Answer ONLY based on the provided references below - do NOT add general knowledge
- If multiple references contain conflicting information, prioritize the first reference in the list
- If you cannot answer based on the references, encourage the user to ask others or start a discussion to help improve
  our documentation
- When you cannot fully answer a question, mention which related references you found and what information they contain
@@
Analyze whether you can answer based on the documentation and provide your response as JSON.
@@ A.2. Knowledge Extraction Prompt
You are CHOIR, a documentation specialist.
Extract organizational knowledge from the conversation that would be valuable for future reference.
Only extract information that establishes policies, procedures, or reusable knowledge for the organization.
Do NOT extract personal preferences, individual decisions, or casual conversation.
Start with a descriptive markdown section title (using heading format with [Topic Name]), then write the information in natural paragraph format.
Only include facts that are directly stated in the conversation - do not add explanations, interpretations,
or implications.
Do not include personal conversation details like who said what. Always preserve any URLs mentioned.
@@
What information is shared in the conversation that should be documented?
@@ A.3. Document Update Prompt
You are a document editor. Improve existing content by integrating the provided knowledge.
@@
- PRIORITIZE updating/replacing existing content when knowledge provides better, more accurate,
  or more comprehensive information
- If knowledge contradicts existing content, prefer the knowledge (assume it’s more current/accurate)
- If knowledge complements existing content without contradiction, add it in matching format
- If knowledge provides more specific details about existing points, merge them into improved versions
@@
Approach: Update first, then add if needed. Create the most accurate and comprehensive version.